OpenStudy (anonymous):

Find the indefinite integral of (y+(2/sin^2y))dy

OpenStudy (jamesj):

Hint: what's the derivative with respect to y of cot(y) ?

OpenStudy (anonymous):

cscx?

OpenStudy (jamesj):

No. But it is related to the 1/sin^2 y in your question

OpenStudy (anonymous):

I still don't get it?

OpenStudy (anonymous):

csc^2x

OpenStudy (jamesj):

\[ \frac{d }{dy} \cot y = - \frac{1}{\sin^2 y} = - \csc^2 y \]

OpenStudy (jamesj):

Hence \[ \int \left( y + \frac{2}{\sin^2 y} \right) \ dy = \int y \ dy + 2 \int \csc^2 y \ dy = ... \]

OpenStudy (anonymous):

So what's the answer?

OpenStudy (jamesj):

You tell me. I've done the first few steps for you. You only have 2 more to do: evaluate each of those integrals above

OpenStudy (anonymous):

y^2/2+2cotx+c

OpenStudy (jamesj):

Double check your signs!

OpenStudy (anonymous):

Negative cotx

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

Yes, thank you!!!